OpenStudy (anonymous):

30% of the students taking algebra also take trigonometry and 50% taking trigonometry also take algebra. if 36 students take both subjects, how many students take only one of the two? please explain! thanks! <3

OpenStudy (anonymous):

no of students taking algebra(x) X x 30/100=36 X=120 out of 120 , 36 take both subjects no of students taking trignometry(y) y x 50 /100 = 36 y=72 out of 72 , 36 take both subjects no of students take only one of the 2 =(120 -36)+(72 -36) =120